OpenStudy (oxblood):

hi guys, some calculus: f(x) has domain [0,infinity] and range [-1,1] g(x) = f^-1(x) f^2(x) = F(x) is inversion possible? F^-1(x) = ?

OpenStudy (anonymous):

doesn't seem likely since if the range of f is [-1,1] then the range of \[f^2\] would be [0,1]

OpenStudy (anonymous):

but that is just a guess. i would be better to think up an example. need a one to one function form (0, infinity) to [-1,1] then maybe we could do it by example

OpenStudy (anonymous):

suppose for example that \[f(a)=-\frac{1}{2}\] then \[f^2(a)=\frac{1}{4}\] but if the range of f is really [-1,1] then \[\frac{1}{4}\] already has a preimage not equal to a in \[(0,\infty)\] so i would say no, function would not be one to one